OpenStudy (anonymous):

solve equation y^2-3y+4=y-1

OpenStudy (anonymous):

\[\sqrt{y^2-3y+4}=y-1\]

OpenStudy (anonymous):

\(y^2-3y+4=y-1\) set equal to zero by subtracting \(y\) and adding \(1\) to both sides and solve \(y^2-4y+5=0\) by using the quadratic formula or completing the square the solutions will be complex (not real)